# references.bib Discovering Dynamic Symbolic Policies with Genetic Programming

### v2

@@ 3.5 Generalising symbolic policies are discovered for industrial application
The average best fitness of the NDE policies has not converged yet after the dedicated number of generations, while being stuck at a poor fitness for the first half of the evolution (Figure 3G, green curve). A potential reason for the lower convergence efficiency of the NDE policies could be that they cannot easily approximate the non-linearity required to stabilise the reactor efficiently, which GP can do more easily, indicated by the faster convergence at a better fitness. The validation fitness of the NDE policies is lower on average than the symbolic policies, and contain a few bad-performing scores (Figure 3G, green dots).
@@
GP-S converges to a lower fitness during evolution than NDE (Figure 3G, blue curve), and the validation fitness shows high consistency among the policies (Figure 3G, blue dots). GP-D obtains a better average best fitness as GP-S (Figure 3G, orange curve). GP-D discovered several policies that outperform GP-S and NDE in terms of validation fitness (Figure 3G, orange dots). However, the high-performing dynamic symbolic policies are again not consistently found, as some policies obtain a validation fitness significantly worse than any static policy.
@@
The best static and dynamic policies are shown in Table 5 (CSTR, both Static and Dynamic). The evolved symbolic policies include non-linear terms such as exponentials and logarithms, but also multiplications of different variables. Such complex non-linear features are more difficult for the NDE to approximate, given the number of neurons in the latent state of the NDE. Increasing the number of neurons would improve its capability to approximate any function, however the optimisation of the network would become less efficient, requiring more computation power. GP can evolve these non-linear functionalities easily, which is yet another advantage of our method.
@@ 3.6 The latent states in a dynamic symbolic policy are transparent
The conducted experiments resulted in many diverse policies with interesting properties. The best validation fitness obtained with each method is presented in Table 4. Generally, the fitness score is the best for the NDE policies, but GP-D is often close and even outperforms NDE for some experiments. The best validation fitness of GP-D is higher than GP-S for every experiment. The corresponding static and dynamic symbolic policies are shown in Table 5. The mathematical expressions in the symbolic policies allow users to interpret a policy’s computations and actions. But in the case of dynamic symbolic policies, the latent states show transparent properties as well. These latent state variables can be plotted during a trial to analyse how the memory capacity is used to solve a problem. We analysed the best policies discovered in various experiments to investigate how these policies solve the task.
@@
Firstly, we investigated the functionality of the best policy evolved to stabilise the SHO with varying parameters settings (Table 5, SHO, Varying parameters, Dynamic). Our interpretation of the symbolic expression of variable in this policy is that it computes the distance between the target and position, adds the velocity and subtracts itself to decrease the influence of the noisy observations. Variable integrates multiplied with a constant, which we interpret as the an estimate of the cumulative error. The readout subtracts and the cube of from the target state to output the policy’s control.
@@
By visualizing the latent states of the policy, we can analyse how the memory capacity aids to generalise to different environment settings. To demonstrate generalisability within a single trial, the policy was subjected to the SHO with parameters that change during the trajectory. When the parameters of the SHO change throughout the trajectory (Figure 4A, top row), the policy should continuously adjust its control signal to stabilise the SHO. The policy manages to keep the state close to the target state by increasing the strength of its negative control signal (Figure 4A, middle row). Variable initially fluctuates until the position is stabilised close to the target, where after it stays close to 0 (Figure 4A, bottom row, blue curve). Variable increases constantly over time to adjust to the changes in the environment, which causes the control force to increase (Figure 4A, bottom row, purple curve).
When the environment instantly changes (Figure 4B, top row), the policy adjusts accordingly and quickly stabilises the SHO in the target state again. Differently from Figure 4A, both latent state variables now react to the changes in the environment. Variable responds with a small bump but stabilises again (Figure 4B, bottom row, blue curve). Initially, variable stabilises once the SHO reaches its target state, but converges to a different value after the environment has changed to adjust the control to appropriate force (Figure 4B, bottom row, purple curve). In both scenarios, the policy robustly adjusts its control to the changes in the environment, and the response of the latent state to the changes is transparent.
@@
When tested on the SHO the latent state of the policy is interpreted relatively easy, because the policy is (almost) fully linear. However, most problems require non-linear solutions. To still offer interpretability of a policy, the latent state should remain transparent when a policy becomes non-linear. To validate the transparency in non-linear policies, the best policy obtained from acrobot swing-up task under full observability is analysed (Table 5, Acrobot, Observation noise, Dynamic). Variable computes the difference between the angles of the links. Variable inputs the angle of the first link in two nested sine functions, multiplied with a scalar. The readout subtracts the cosine of from . Our interpretation of this policy is that it applies force in the direction of which the first link is compared to the centre point to build momentum on the first link, and increases the force when the two links are not closely aligned to add extra momentum to the second link, resembling a bang-bang controller.
@@
The mathematical expressions of the non-linear policy to solve the acrobot are less straightforward to interpret than the linear policy that stabilises the SHO. Nonetheless, analysing the latent state of the non-linear policy could help with understanding the functioning of this policy. The policy accomplishes the swing up of the acrobot in a short time (Figure 4C, top row). The control force applied by the policy shows expected behaviour, switching between -1 and 1 to build momentum, and when the second link is close to the threshold the control force remains constant for longer (Figure 4C, middle row). The latents show similar patterns, although variable fluctuates with a higher amplitude (Figure 4C, bottom row). The latent states are not as expressive as in the policy that was applied to the SHO, and it is difficult to understand the individual contribution of each variable. Nonetheless, there is still more transparency in these dynamics than would be the case in a black-box policy.
@@
The previous experiments showed that the dynamic symbolic policies are transparent, but the problems of interest are simple control problems. For application purposes, it is more relevant if the policies remain interpretable in an industrial setting, such as the CSTR. The best policy discovered in the experiment conducted on the CSTR is tested in a single trial to analyse the control and memory of the policy (Table 5 CSTR, Dynamic). We interpret the state equation of variable as a scaling factor that converges to 2.92 during the trajectory. The state equation of variable computes the error between the temperature and setpoint, and adds a self-connection to make the error estimate less affected by noise observations. The error estimate is scaled by the sum of the control, itself and a constant. The readout scales with the squared of , and adds a multiplication of with the logarithm of the setpoint temperature. The control therefore reacts to differences between the setpoint and current temperature, which is scaled by a factor that increases during the trajectory.
@@
To increase the number of transparent responses in the trial, the setpoint temperature in the CSTR is changed several times during the trajectory. Therefore the policy has to adjust its control signal to change the temperature accordingly. The policy efficiently stabilises the temperature around each setpoint temperature (Figure 4D, top row). The control signal of the policy heavily changes when a new setpoint temperature is required, even remaining close to zero when the temperature has to increase. The concentration of the reactant in the main reactor goes to zero quickly and only slightly increases when the control input changes (Figure 4D, middle row). Variable converges to 2.92 as expected (Figure 4D, bottom row, blue curve), while shows a large response when the setpoint temperature changes. Together, the latent variables allow the policy to change its control output accordingly.
@@
All these example trials demonstrate that the dynamic symbolic policies are not only interpretable, but also offer transparency in the latent states. The latents vary in their expressiveness, but overall analysing the latent state aids with understanding the functionality of a policy. The memory capacity already showed to be beneficial for increasing the robustness of a policy, however it also further improves the interpretability of symbolic policies.
@@ 4 Discussion
In this paper, we evolved white-box symbolic policies with genetic programming (GP) to solve control problems. The goal was to produce high-performing policies with interpretable dynamics. We considered two types of symbolic policies; static policies that are memory-less, and dynamic policies that have memory. The static policies include an expression for every control signal required, but the dynamic policies consist of a latent state and a readout function, for which individual trees were optimised with GP. We showed that GP was able to successfully evolve both static and dynamic symbolic policies for solving a variety of classic control problems. The dynamic policies allowed to perform successful control under noisy and partial state observations, as well as different and changing environments. We demonstrated the results in linear and a non-linear environment, and finally for an industrial environment. Overall, the results show that our method can optimise high-performing static and dynamic symbolic policies.
Currently, most approaches for discovering symbolic expressions with GP optimise single trees or apply tricks to evolve multiple trees independently. In the context of control, this would mean that only single dimensional control policies could be learned, resulting in purely static policies. By adding a latent state to the dynamic policies, the policies obtain memory capacity, by which we could improve their performance in dynamic environments. The latent state comes with the cost of requiring GP to optimise multiple trees together. In the experiments where the full state of the environment was observed, both the dynamic and static symbolic policies find policies that perform successfully. The search space of the dynamic policies is substantially larger, as is observed from the poor initial fitness compared to the initial fitness of the static policies and the fewer generations the static approach needs to converge. The validation fitness of the best static policies evolved in each run shows that these policies consistently converge to a similar expression and score. The validation fitness of the dynamic policies displays a larger spread, indicating that the evolution does not consistently result in policies of similar quality. The inconsistency in evolution can also be explained by the significantly larger search space of the dynamic policies.
The larger search space and consequently less consistent optimisation is a disadvantage of the addition of a latent state to the policies. However, the experiments with partial state observability show that the latent state also significantly improves the robustness of the policies. Static policies are unable to solve the tasks when the observations are incomplete, while the dynamic policies achieve similar performance as in the experiments with observation noise only. In environments with full state observability, our method discovers multiple policies that are better than any static policy, although these policies are not found consistently. Overall, our method evolves dynamic symbolic policies efficiently given the large search space, and the resulting policies are generally better than the static policies.
One way to improve the consistency and convergence rate of the evolution of dynamic symbolic policies would be to increase computational power. The hyperparameters of GP that relate to the computation power are the population size and number of generations. These hyperparameters were chosen to be similar to values commonly used in literature to demonstrate that our method evolves these policies relatively efficiently. Nonetheless the convergence rate of our method could be improved by increasing these hyperparameters, which should increase the chance of evolving policies that compete with the black-box baseline.
The black-box method that was used for comparison is a neural differential equation optimised with covariance-matrix adaptation evolution strategies, as this approach has been proven to be powerful alternative to gradient descent optimisation in reinforcement learning settings [salimans2017evolution]. The results demonstrate that symbolic policies are evolved that generally obtain comparable performance as the black-box baseline. The black-box baseline does have more consistent performance, but the white-box policies have several other advantages over the black-box policies. The main advantage of the symbolic policies is that the expressions of the white-box policies are interpretable and can be analysed to gain new insights. Additionally, the behaviour of the latent states is transparent, as demonstrated in Figure 4. This focus on interpretable policies relates to the field of scientific machine learning [rackauckas2020universal]. In scientific machine learning, knowledge and insights about data is sought with data-driven approaches. Since the latent and readout functions in our method were not directed to learn specific properties, the policies could discover different interesting features from scratch in every evolutionary run. Examples of learned properties are error estimates, reliable estimates of noisy observations, features from the environment and bang-bang control (Table 5).
While results are promising, there is room for future improvement of our method for evolving dynamic symbolic policies, both in terms of performance of the discovered policies and the efficiency of the search algorithm. GP in general is inefficient for optimisation of constants, because the search space is extremely large. Hybrid GP is an extension that combines standard GP with numerical optimisation, such as gradient-based methods or evolution strategies [topchy2001faster]. Focused optimisation of the constants improves the performance of the policy, but it could also save time during the search process of GP. Hybrid GP either optimises the constants of the final policy, or constant optimisation is integrated in the evolution progress itself, as was done in [cao2000evolutionary].
Another improvement would be to vary the number of latent states through evolution as well, similar to the NEAT algorithm [stanley2002evolving]. With GP the size of the latent state in a policy can be increased when more memory capacity is beneficial to solve a problem. Similarly the number of latents can be decreased when not all memory capacity is necessary to solve a problem. The search space grows even more with this implementation, however the policies will consist only of relevant trees. Besides the described options there are many improvements for more efficient optimisation in GP, such as semantic reproduction [vanneschi2014survey], population diversity control [burke2004diversity] and self-adaptation of hyperparameters [angeline1996two].
@@
Note that we chose to learn symbolic expressions for both the latent state and readout functions. To further increase the interpretability in the discovered policies, the readout could be learned as a vector that is multiplied with the latent state to produce a control signal. This way, all non-linear functionality is evolved in the latent state, and the relation between the latent state and control signal becomes easier to understand. The robustness of evolution also improves, as less trees have to be optimised with GP, although another method for optimising the readout vector should be introduced.
@@
Lastly, a major limitation of GP, which holds especially for the dynamic policies, is the huge search space. Our method showed to be scalable when the environment required two control outputs. Nonetheless, it is important to test our method on problems that receive higher-dimensional control and more observations, such as robot control. Future work would be to extend our research to high-dimensional problems, and to analyse how the GP algorithm should be improved to deal with the increased search space efficiently.
We conclude that the use of genetic programming is a viable route towards learning interpretable, efficient and general dynamic symbolic policies for control. We hope that this adds an important tool to the practitioners’ toolbox in the endeavour to create trustworthy AI systems.
@@ Appendix A Genetic programming
| SHO | Acrobot | CSTR | |||||
| Hyperparameter | Exp 1 | Exp 2 | Exp 3 | Exp 1 | Exp 2 | Exp 3 | Exp 1 |
| Population Size | 500 | 500 | 1000 | 500 | 500 | 1000 | 1000 |
| Number of Generations | 30 | 50 | 100 | 50 | 50 | 50 | 100 |
| Latent state size | 2 | 2 | 2 | 2 | 2 | 3 | 2 |
| Function set | , power | power, sin, cos | power, exp, log | ||||
@@ Appendix B Acrobot parameters
| rad | Angle of a link | ||
| rad/s | Angular velocity of a link | ||
| m | Length of a link | 1.0 | |
| kg | Mass of a link | 1.0 | |
| m | Position of the center of mass of a link | 0.5 | |
| kg | Moment of inertia of a link | 1.0 | |
| m/ | Gravity | 9.81 |
@@ Appendix C Continuous stirred tank reactor parameters
| gmol/liter | Reactant concentration | ||
| K | Reactor temperature | ||
| K | Cooling jacket temperature | ||
| K | Setpoint temperature | {400, 500} | |
| liters/min | Feed flowrate of reactant into reactor | {75, 125} | |
| liters/min | Feed flowrate of coolant into cooling jacket | Controlled parameter | |
| gmol/liters | Feed concentration of reactant into reactor | 1 | |
| K | Feed temperature of reactant into reactor | {300, 350} | |
| K | Feed temperature of coolant into cooling jacket | {250, 300} | |
| liters | Volume of reactor | {75, 150} | |
| liters | Volume of cooling jacket | {10, 30} | |
| J/gmol | Enthalpy of reaction | {-55000, -45000} | |
| g/liters | Density | 1000 | |
| J/g/K | Heat capacity | {0.2, 0.35} | |
| J/min/K | Heat transfer coefficient | {25000, 75000} | |
| 1/min | Arrhenius pre-exponential | ||
| J/gmol | Activation energy | 72750 | |
| J/gmol/K | Gas constant | 8.314 | |
| 1/min | Reaction rate | Function of temperature |
Appendix D Discovered policies
Table 4 shows the best validation fitness per experiment obtained by static symbolic policies (GP-S), dynamic symbolic policies (GP-D), and the two baselines; random search (RS) and the black-box neural different equations (NDE). GP-D outperforms GP-S in every experiment, and sometimes even outperforms NDE. The corresponding best static and dynamic symbolic policies discovered in each experiment are presented in Table 5. The static policies contain as many expressions as the environment has inputs. The dynamic policies contain another set of expressions for the state equations of the latent state.
| Experiment | RS | NDE | GP-S | GP-D |
| Harmonic oscillator (Observation noise) | -56.11 | -48.85 | -57.28 | -49.66 |
| Harmonic oscillator (Partial state observability) | - | -72.76 | -903.44 | -63.35 |
| Harmonic oscillator (Varying parameters) | - | -47.76 | -110.25 | -45.56 |
| Acrobot (Observation noise) | -123.64 | -98.45 | -115.31 | -112.82 |
| Acrobot (Partial state observability) | - | -101.47 | -250.0 | -110.23 |
| Acrobot (Two control inputs) | - | -97.02 | -103.43 | -100.69 |
| CSTR | - | -1017.72 | -1059.31 | -845.33 |
| Experiment | Static | Dynamic |
| SHO | ||
| Observation noise | ||
| SHO | ||
| Partial state observability | ||
| SHO | ||
| Varying parameters | ||
| Acrobot | ||
| Observation noise | ||
| Acrobot | No successful policies | |
| Partial state observability | ||
| Acrobot | ||
| Two control inputs | ||
| CSTR | ||

### v3

@@ 3.5 Generalising symbolic policies are discovered for an industrial application
The NDE has not converged after the dedicated number of generations, while being stuck at a poor fitness for the first half of the evolution (green curve). A potential reason for the limited convergence of the NDE could be that they cannot easily approximate the non-linearity required to stabilise the reactor efficiently. The validation fitness of the NDE policies is lower on average than the symbolic policies, and there are a few bad-performing policies (green dots).
@@
GP-S converges to a higher fitness than NDE (blue curve), and the validation fitness shows high consistency among the policies (blue dots). GP-D obtains a better fitness than GP-S during evolution (orange curve). GP-D discovered several policies that outperform GP-S and NDE in terms of validation fitness (orange dots). However, the high-performing dynamic symbolic policies are again not consistently found, as some policies obtain a validation fitness significantly worse than any static policy.
@@
The best static and dynamic policies are shown in Table 1 (CSTR, both Static and Dynamic). The evolved symbolic policies include non-linear terms such as logarithms and multiplications of different variables. Such complex non-linear features are more difficult to approximate for the NDE, given the number of neurons in the latent state of the NDE. Increasing the number of neurons would improve its capability to approximate any function, however the optimisation of the network would become less efficient, requiring more computation. GP evolves these non-linear functionalities easily, which highlights another advantage of our method.
| Environment | Experiment | Static policy | Dynamic policy |
| SHO | Observation noise | ||
| SHO | Partial state observability | ||
| SHO | Varying parameters | ||
| Acrobot | Observation noise | ||
| Acrobot | Partial state observability | No successful policies | |
| Acrobot | Two control inputs | ||
| CSTR | Standard |
@@ 3.6 The latent states in a dynamic symbolic policy are transparent
Across the experiments, GP was able to find successful dynamic symbolic policies. The best validation fitness obtained with each method is presented in Table 4h. GP-D is generally competitive with black-box NDEs, and outperforms the NDE in 3 out of 7 experiments. The best validation fitness of GP-D is higher than GP-S for every experiment. The dynamic symbolic policies compete with the black-box baseline, while consisting of interpretable mathematical expressions. The best static and dynamic symbolic policies are displayed in Table 1. Visualising the dynamic symbolic policies reveals that the latent states exhibits transparent properties that allow us to investigate how the memory is used to solve control tasks.
@@
First, we investigated the functionality of the best policy evolved to stabilise the SHO with varying parameters settings (Table 1, SHO, Varying parameters, Dynamic). Our interpretation of the symbolic expression of variable is that it computes the distance between the target and position, adds the velocity and subtracts itself to decrease the influence of the noisy observations. Variable integrates multiplied with a constant, which we interpret as the estimate of the cumulative error. The readout subtracts and the cube of from the target state to output the policy’s control.
@@
To demonstrate the generalisability of this policy within a single trial, we change the parameters of the SHO during the experiment (Fig. 5a, top row). The policy manages to keep the state close to the target state by increasing the strength of its negative control signal (Fig. 5a, middle row). Variable initially fluctuates until the position is stabilised close to the target, where after it stays close to the origin (Fig. 5a, bottom row, red curve). Variable first decreases, but increases during the trajectory to adjust to the changes in the environment, which causes the control force to increase (Fig. 5a, bottom row, blue curve).
@@
The latent state should remain transparent when a policy becomes non-linear to still offer interpretability in more complex tasks than the SHO. To demonstrate the transparency in non-linear policies, the best policy obtained from acrobot swing-up task with observation noise is analysed (Table 1, Acrobot, Observation noise, Dynamic). Variable computes the difference between the angles of the two links. Variable inputs the angle of the first link in two nested sine functions, multiplied with a scalar. The readout subtracts the cosine of from . Our interpretation of this policy is that it applies force in the direction of the first link compared to the centre point to build momentum, and increases the force when the two links are not closely aligned to add extra momentum to the second link, resembling a bang-bang controller.
@@
The mathematical expressions of the non-linear policy are less straightforward to interpret than the linear policy. Analysing the latent state of this policy helps with understanding the functioning of this policy. The policy quickly accomplishes the swing up of the acrobot (Fig. 5b, top row). The control force applied by the policy shows expected behaviour, switching between -1 and 1 to build momentum, and remains constant longer when the second link is close to the threshold height (Fig. 5b, middle row). The latent state shows similar behaviour, although variable fluctuates with a higher amplitude (Fig. 5b, bottom row). The latent states are not as expressive as in the policy in Fig. 5a, and it is difficult to understand the individual contribution of each variable. Nonetheless, there is still more transparency in these dynamics than in a black-box policy.
@@
For application purposes, it is also relevant that policies remain interpretable in an industrial setting, such as the CSTR. The best policy discovered for stabilising the CSTR is analysed for transparency (Table 1 CSTR, Dynamic). We interpret the state equation of variable as a scaling factor that converges to 2.92 during the trajectory. The state equation of variable computes the error between the temperature and setpoint, and adds a self-connection to improve robustness to noise. The error estimate is scaled by the sum of the control, and a constant. The readout scales with the squared of , and adds a multiplication of with the logarithm of the setpoint temperature. The control therefore reacts to differences between the setpoint and current temperature, scaled increasingly during the trajectory.
@@
To increase the number of responses in the trial, the setpoint temperature is changed several times during the trajectory. The policy efficiently stabilises the temperature at each setpoint temperature (Fig. 5c, top row). The control signal reacts when the setpoint temperature changes, and remains close to zero when the temperature has to increase. The concentration of the reactant in the main reactor goes to zero quickly and only slightly increases when the control input changes (Fig. 5c, middle row). Variable converges to 2.92 as expected (Fig. 5c, bottom row, red curve), while shows large responses when the setpoint temperature changes (Fig. 5c, bottom row, blue curve). Together, the latent variables allow the policy to change its control output accordingly.
@@
These example trials demonstrate that the dynamic symbolic policies are not only interpretable, but also offer transparency in the latent states. The latent state variables vary in their level of expressiveness, but overall analysing the latent state aids with understanding the functionality of a policy. The latent state showed to be beneficial for the robustness of a policy, however it also further improves the interpretability of symbolic policies.
@@ 4 Discussion
In this paper, we discovered interpretable and high-performing white-box control systems to solve control problems. The conventional approach is to find static symbolic policies, but these are difficult to extend to real-world challenges like partial observability and volatility. To improve the robustness of control systems, we evolved dynamic symbolic policies with memory with genetic programming (GP). The dynamic policies successfully solved control tasks with noisy and partial state observations, where the static policies failed to solve the task. Furthermore, dynamic symbolic policies generalise better to tasks with varying environment settings than static policies. These results show the benefits of the memory in the dynamic symbolic policies, and these policies can still be evolved efficiently with GP in terms of generations and population size. The symbolic policies were compared with a neural differential equation optimised with covariance-matrix adaptation evolution strategies as a proven black-box baseline [salimans2017evolution]. The results demonstrate that dynamic symbolic policies are evolved that generally obtain comparable performance as the black-box baseline, although competitive symbolic policies are not consistently found across evolutionary runs.
@@
Besides adequately solving control tasks, the dynamic symbolic policies have other advantages, such as that they consist of interpretable expressions and display transparent dynamics, as demonstrated in Fig. 5. These features help understanding the functioning of control systems, therefore contributing to identification of undesired behaviour. Additionally, the interpretability allows us to learn new approaches to solve tasks from the policies. To further increase the interpretability in the dynamic symbolic policies, the readout could be learned as a numerical vector that maps the latent state to a control signal. This way, all non-linear functionality is evolved in the latent state, and the relation between the latent state and control signal becomes easier to understand, which is currently not always trivial (Table 1). The consistency of evolution also improves, as less trees have to be optimised with GP, although another method for optimising the readout vector should be introduced.
The major limitation of evolving dynamic symbolic policies with GP, is the huge search space that has to be traversed, which hinders consistent evolution of high-performing policies. The results showed that GP converges faster and more consistently when static policies are evolved under standard environment circumstances. The search space will grow even larger when the control problem contains higher-dimensional observation and control spaces. Our method demonstrated to be scalable when the environment required two control outputs (Fig. 4f), but this may not hold for more complex problems. Better policies can be evolved with GP when a higher population size and number of generations are selected, which also increases the computation requirements. However, previous research has focused extensively on improving the optimisation efficiency of GP, by integrating features such as optimisation of constants [topchy2001faster], semantic reproduction [vanneschi2014survey], population diversity control [burke2004diversity] and self-adaptation of hyperparameters [angeline1996two].
Besides introducing the concept of dynamic symbolic policies, this research also resulted in a general GP framework that is not limited to optimisation of dynamic symbolic policy. First of all, other structures of symbolic policies can be configured, with the possibility to explicitly define different types of trees with unique inputs and functionality in the policy. Furthermore, the framework could also be used problems outside control, for example symbolic regression [bongard2007automated], discovering learning rules [jordan2021evolving] or other directions within scientific discovery [wang2023scientific, vinuesa2024opportunities, rackauckas2020universal].
@@
We conclude that the use of genetic programming is a viable route towards learning interpretable, efficient and robust dynamic symbolic policies for solving control tasks. We hope that this adds an important tool to the practitioners’ toolbox in the endeavour to create trustworthy AI systems.
@@ Appendix A Genetic programming
| SHO | Acrobot | CSTR | ||||
| Hyperparameter | Exp 1 | Exp 2 | Exp 3 | Exp 1 | Exp 2 | Exp 1 |
| Population Size | 500 | 500 | 1000 | 500 | 500 | 1000 |
| Number of Generations | 30 | 50 | 100 | 50 | 50 | 100 |
| Latent state size | 2 | 2 | 2 | 2 | 2 | 2 |
| Function set | , power | power, sin, cos | power, exp, log | |||
@@ Appendix B Acrobot parameters
[…]
@@ Appendix C Continuous stirred tank reactor parameters
[…]